Question 12 Find an equation of the form y = a* + k for the graph sketched below: -5 -4 -3 -2 - f(x) =
Answer #1

Given y = ah+k at a=0 g = 4 4 = dotky 4 =1tk. = K = 3 ...y - a2 +3 at x = -1 y = 6 6= a +3 a-t=3 as 7. () + Ang